CONTROL OF WATER CLUSTER GROWTH BY Ne INSERTION IN He NANODROPLETS

\underline{BRIAN E. APPLEGATE}, AND ROGER E. MILLER, University of North Carolina, Dept. of Chemistry, Chapel Hill, NC 27599; CHRISTIAN J. BURNHAM AND SOTIRIS S. XANTHEAS, Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, 906 Battelle Boulevard, MS K8-91, Richland, WA 99352.

We have shown that the growth of the (H_{2}O)_{n} clusters in He nanodroplets proceeds via the sequential addition of water molecules to the preformed (H_{2}O)_{n−1} rings, follwed by ring insertion. However, if we add a small number of Ne atoms to the He nanodroplet before commencing the water cluster growth process, a polar water complex is formed, presumably because the Ne stifles this ring insertion mechanism. The spectroscopy and corresponding ab initio calculations of this polar complex will be discussed.